# Prior-based Noisy Text Data Filtering: Fast and Strong Alternative For Perplexity

### v2

@@ 4.2 Experiment on symbolic language corpus
Table 3: Performance of the baselines pre-trained on Pile-github across 6 symbolic problem solving benchmarks
Time
Average
normalized
accuracy
BIG-bench
cs algorithms
BIG-bench
dyck languages
BIG-bench
operators
BIG-bench
elementary
math QA
GSM8K SVAMP
Large (1.5B) model
@@ Appendix A Related works
In this section, we first review two streams (e.g., rule-based and model-based) of previous works on web text data filtering for the pretraining of LLMs. Next, we more closely describe those sharing conceptual similarities with our proposed method.
@@
Raw web-scraped data often contains a substantial amount of low-quality content, including documents with only space or machine-generated spam [20]. Because such noise can be detected with only very simple rules, the earliest filtering methods were rule-based. [32] introduced the following rules, which have subsequently been adopted in a similar form by most later works [22, 30, 10, 31, 5]: (1) Retain lines only with a terminal punctuation mark. (2) Discard pages with fewer than 3 sentences. (3) Keep lines with at least 5 words. (4) Remove pages with words from the bad-word list. (5) Removed lines containing the word “Javascript” and “lorem ipsum”.
Because of its simplicity, such heuristic methods often fail to apply fine-grained filtering and risk discarding semantically valuable content inadvertently [13]. Nevertheless, because rule-based methods are extremely lightweight and require minimal computational resources, they are commonly applied at the web crawling stage [7, 6]. Consequently, most web datasets already have such methods applied by default [30, 39].
@@
(1) Without reference: A representative method is computing the perplexity of an LM on the text and filtering out outliers according to this measure. Another approach employs a linear classification method based on the embeddings of an LLM [22]. AskLLM [36] operates by presenting data points to the LLM and utilizing its reasoning capability to judge whether they constitute noisy data. EL2N [29] ranks samples based on the L2 distance between a model’s prediction and the ground truth, thereby identifying data points that are more important for learning. Similarly, memorization-based methods [8] assess how well a model memorizes token sequences within a document. PageRank score is a method of filtering documents based on how likely they are to be hyperlinked to other documents [22]. Semantic Deduplication (semDedup) embeds each point in a dataset, clusters data points together, and removes data points within clusters that are too similar [1].
(2) With reference: A well-known method is DSIR [44], which constructs embeddings using either n-grams or FastText embeddings on a curated dataset, and evaluates the similarity of samples in the raw dataset to these reference embeddings. DSIR is proposed to advance FastText [18]; it trains the model to classify a curated dataset as either high quality or low quality.
@@
PPL-based and DSIR are the most advanced
A review of the representative literature [22, 4, 23, 44] shows that PPL-based and DSIR are the most advanced filtering methods to date. First, according to these studies, model-based methods significantly outperform rule-based ones. Furthermore, [22] found that PPL-based outperforms or is comparable to other model-based methods (e.g., PageRank, semDedup, LLM-based classifiers, and AskLLM) and is comparable with FastText. While [22] selected low PPL data, [4] found that selecting median PPL data achieves significantly better performance. Also, [44] proposed DSIR as an improved version of fastText, with higher performance. In conclusion, we get PPL-based and DSIR as the most advanced methods to compare as baselines.
A.1 Methods with conceptual relation
A.1.1 DSIR
DSIR [44] assumes that a well-curated reference dataset consisting of high-quality, well-formed text is available (Wikipedia and Bookcorpus is used in the original work). The method is to evaluate the similarity of sample $\mathtt{d}$} in the raw dataset to this reference corpus, and uses it as the filtering criterion.
According to [44], the process for estimating this similarity proceeds as follows. Given a corpus $D$, each document ${\mathtt{d}}\in D$ is sliced into a sequence of $n$-grams. For example, if the text input is “Alice is eating”, it forms the list [Alice, is, eating, Alice is, is eating]. These $n$-grams are then mapped to hash indices, which are subsequently grouped into $m$ hash buckets (with $m=10000$). The resulting hash frequencies form an $m$-dimensional categorical distribution vector $\gamma\in\mathbb{R}^{m}$, referred to as the feature distribution $P$. Separate feature distributions $P_{raw}$ and $P_{ref}$ are computed for the reference dataset and the raw dataset, respectively (each denoted as $q$ and $p$ in the original paper).
@@ A.1.1 DSIR
A.1.2 SCDP
SCDP (Swift Cross-Data Pruning) [26] is a method that selects data based on the multivariate median of TF-IDF (term frequency and inverse document frequency) representations. This method selects data that is most similar to the dominant topic frequently covered in the corpus.
@@ A.1.2 SCDP
A.1.3 FastText
FastText [18] is a model-based filtering method. To train the model, web-crawled data are assigned as “low-quality,” while manually curated target data are assigned as “high-quality”. And the classifier model is trained to infer the probability of a datapoint belonging to a high-quality set. [44] explains that DSIR is an improved version of FastText, outperforming in performance. Like DSIR, a key limitation of FastText is that filtering quality is constrained by the accuracy of manual curation, which also increases rigidity and complexity due to the required human labor. Moreover, because web-crawled data must be additionally tokenized, the computational time cost increases.
We provide additional analysis on the FastText model in the same setting as DCLM [22], assigning web-crawled data (RefinedWeb [30], which is based on Common Crawl) as the “low-quality” set, while ELI5 and OH-2.5 as the “high-quality” set.
@@ A.1.3 FastText
To avoid the cost of human labeling, DCLM chose to label RefinedWeb [2] as the low-quality set, while labeling OH-2.5 and ELI5 as the high-quality set. During inference, the model predicts the probability (0 100%) that a given text belongs to the high-quality set. However, RefinedWeb also contains a considerable proportion of well-formed documents. As a result, many samples that FastText classifies with 0% probability of belonging to the high-quality set are in fact well-formed and informative texts (Case1), often resembling those found in RefinedWeb.
@@
Appendix B Additional analysis
Table 4: Performance and time cost (for filtering) of the baselines pre-trained on Dolma.
Criteria Tokenizer Threshold Source
Average
normalized
accuracy
World
knowledge
Commonsense
reasoning
Language
understanding
Symbolic
problem
solving
Reading
compre-
hension
Large (1.5B) model
prior-based mean GPT-2 50% Dolma 8.50 9.12 10.25 7.45 11.38 4.28
" stds GPT-2 " " 8.70 7.28 10.57 9.34 12.40 3.89
" mean + stds GPT-2 " " 9.20 9.53 11.27 10.31 11.13 3.79
" " LLaMA-3 " " 9.39 9.54 11.16 10.78 11.86 3.64
" " T5-small " " 8.11 8.59 7.43 7.95 12.32 4.22
" " GPT-2 elbow (81%) " 8.79 10.04 7.67 8.97 13.04 4.21
" " GPT-2 50% Dolma + Github 9.48 11.47 10.83 8.97 12.22 3.78
We present detailed analyses of several important issues. The related ablation experiments were conducted on a large model using Dolma (Table 4).
B.1 Preserving minority data

### v4

@@ 4.2 Experiment on symbolic language corpus
Table 3: Performance of the baselines pre-trained on Pile-github across 6 symbolic problem-solving benchmarks. CS: BIG-bench cs algorithms, Dyck: BIG-bench dyck languages, Op: BIG-bench operators, Ma: BIG-bench elementary math QA, GSM: GSM8k, SVA: SVAMP
Time Avg CS Dyck Op Ma GSM SVA
1.5B model
@@ 5 Conclusion and limitation
Acknowledgement
This research was supported by Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government(MSIT), under AI Technology Development for Commonsense Extraction, Reasoning, and Inference from Heterogeneous Data(2022-0-00077, RS-2022-II220077, 50%), ITRC(Information Technology Research Center) support program(IITP-2025-RS-2024-00437102, 30%), Artificial Intelligence Graduate School Program (Yonsei University)(RS-2020-II201361, 20%). Jinyoung Yeo and Jaehyung Kim are the co-corresponding authors.
@@ Appendix A Related works
In this section, we first review two streams (e.g., rule-based and model-based) of previous works on web text data filtering for the pretraining of LLMs. Next, we explain that the baselines employed in our experimental setup cover all categories of existing baselines. Finally, we provide a more in-depth explanation of previous methods that share conceptual similarities with our proposed method.
A.1 Major streams of prior work
@@
Raw web-scraped data often contains a substantial amount of low-quality content, including documents with only space or machine-generated spam (Kreutzer et al., 2022). Because such noise can be detected with only very simple rules, the earliest filtering methods were rule-based. Raffel et al. (2020) introduced the following rules, which have subsequently been adopted in a similar form by most later works (Li et al., 2024a; Penedo et al., 2023; Chen et al., 2021; Rae et al., 2021; Bane et al., 2022): (1) Retain lines only with a terminal punctuation mark. (2) Discard pages with fewer than 3 sentences. (3) Keep lines with at least 5 words. (4) Remove pages with words from the bad-word list. (5) Removed lines containing the word “Javascript” and “lorem ipsum”.
Because of its simplicity, such heuristic methods often fail to apply fine-grained filtering and risk discarding semantically valuable content inadvertently (Dodge et al., 2021). Nevertheless, because rule-based methods are extremely lightweight and require minimal computational resources, they are commonly applied at the web crawling stage (Bevendorff et al., 2018; Barbaresi, 2021). Consequently, most web datasets already have such methods applied by default (Penedo et al., 2023; Soldaini et al., 2024).
@@
(1) Without reference: A representative method is computing the perplexity of an LM on the text and filtering out outliers according to this measure. Another approach employs a linear classification method based on the embeddings of an LLM (Li et al., 2024a). AskLLM (Sachdeva et al., 2024) operates by presenting data points to the LLM and utilizing its reasoning capability to judge whether they constitute noisy data. EL2N (Paul et al., 2023) ranks samples based on the L2 distance between a model’s prediction and the ground truth, thereby identifying data points that are more important for learning. Similarly, memorization-based methods (Biderman et al., 2023) assess how well a model memorizes token sequences within a document. PageRank score is a method of filtering documents based on how likely they are to be hyperlinked to other documents (Li et al., 2024a). Semantic Deduplication (semDedup) embeds each point in a dataset, clusters data points together, and removes data points within clusters that are too similar (Abbas et al., 2023).
(2) With reference: A well-known method is DSIR (Xie et al., 2023b), which constructs embeddings using either n-grams or FastText embeddings on a curated dataset, and evaluates the similarity of samples in the raw dataset to these reference embeddings. DSIR is proposed to advance FastText (Joulin et al., 2016); it trains the model to classify a curated dataset as either high quality or low quality.
@@
A.2 Explanation on our baseline settings
As shown in Table 2, the baselines used in our main experiments include the vanilla Dolma dataset, FastText, DSIR, and PPL-based filtering. We explain how these methods comprehensively cover the major streams of baselines described above.
(1) vanilla Dolma (noted as “no-filter” in the paper)
Dolma (Soldaini et al., 2024) corpus itself already incorporates multiple filtering stages—including heuristic filtering, quality filtering, and even model-based filtering.
  • Heuristic rule: adopted from rules used in C4, Gopher.
  • Language filter (classifier-based) : Using fastText to select English data.
  • Content filter (classifier-based) : Using fastText trained on the Jigsaw Toxic Comment dataset to filter out toxic documents
  • Deduplication (rule-based) : duplication regarding an exact match of URL, document, or paragraph.
Our method achieves substantial performance gains on top of this. Moreover, these filters are usually computationally lightweight and are often applied already during the web-crawling stage (Li et al., 2024a; Soldaini et al., 2024). Therefore, they are compatible with our method by default.
(2) DSIR
DSIR operates on n-gram embedding vectors trained on a curated dataset. This work (Xie et al., 2023b) also compares a classifier-based embedding option (FastText). So this baseline covers both n-gram and classifier-based methods.
(3) PPL-based
The (Li et al., 2024a) paper compares a wide range of methods, and demonstrates that PPL-based approach outperforms well-known model-based methods (AskLLM, classifier on BGE features, PageRank, SemDedup) and heuristic filters, or is comparable to FastText and top-k average logits. Therefore, we selected the PPL-based method as the baseline representing other classifier-based methods.
(4) FastText
To directly compare against classifier-based approaches, we additionally provide the comparison with FastText (trained under the same dataset as DSIR). Its performance was similar but slightly lower than DSIR, which is consistent with the findings reported in (Xie et al., 2023b). The limitations of curation-based methods are also discussed in §A.3.3.
Conclusion
In summary, our experiments cover various types of baselines, including heuristic rule-based, n-gram-based, and classifier-based methods. In addition, we provide an extra classifier-based baseline experiment for the sake of clarity.
A.3 Methods with conceptual relation
A.3.1 DSIR
DSIR (Xie et al., 2023b) assumes that a well-curated reference dataset consisting of high-quality, well-formed text is available (Wikipedia and Bookcorpus is used in the original work). The method is to evaluate the similarity of sample $\mathtt{d}$} in the raw dataset to this reference corpus, and uses it as the filtering criterion.
According to Xie et al. (2023b), the process for estimating this similarity proceeds as follows. Given a corpus $D$, each document ${\mathtt{d}}\in D$ is sliced into a sequence of $n$-grams. For example, if the text input is “Alice is eating”, it forms the list [Alice, is, eating, Alice is, is eating]. These $n$-grams are then mapped to hash indices, which are subsequently grouped into $m$ hash buckets (with $m=10000$). The resulting hash frequencies form an $m$-dimensional categorical distribution vector $\gamma\in\mathbb{R}^{m}$, referred to as the feature distribution $P$. Separate feature distributions $P_{raw}$ and $P_{ref}$ are computed for the reference dataset and the raw dataset, respectively (each denoted as $q$ and $p$ in the original paper).
@@ A.3.1 DSIR
A.3.2 SCDP
SCDP (Swift Cross-Data Pruning) (Nguyen and He, 2025) is a method that selects data based on the multivariate median of TF-IDF (term frequency and inverse document frequency) representations. This method selects data that is most similar to the dominant topic frequently covered in the corpus.
@@ A.3.2 SCDP
A.3.3 FastText
FastText (Joulin et al., 2016) is a model-based filtering method. To train the model, web-crawled data are assigned as “low-quality,” while manually curated target data are assigned as “high-quality”. And the classifier model is trained to infer the probability of a datapoint belonging to a high-quality set. (Xie et al., 2023b) explains that DSIR is an improved version of FastText, outperforming in performance. Like DSIR, a key limitation of FastText is that filtering quality is constrained by the accuracy of manual curation, which also increases rigidity and complexity due to the required human labor. Moreover, because web-crawled data must be additionally tokenized, the computational time cost increases.
We provide additional analysis on the FastText model in the same setting as DCLM (Li et al., 2024a), assigning web-crawled data (RefinedWeb (Penedo et al., 2023), which is based on Common Crawl) as the “low-quality” set, while ELI5 and OH-2.5 as the “high-quality” set.
@@ A.3.3 FastText
To avoid the cost of human labeling, DCLM chose to label RefinedWeb (Penedo et al., 2023) as the low-quality set, while labeling OH-2.5 and ELI5 as the high-quality set. During inference, the model predicts the probability (0 100%) that a given text belongs to the high-quality set. However, RefinedWeb also contains a considerable proportion of well-formed documents. As a result, many samples that FastText classifies with 0% probability of belonging to the high-quality set are in fact well-formed and informative texts (Case1), often resembling those found in RefinedWeb.
@@
Appendix B Theoretical derivation and boundary condition analysis
In this section, we show that (1) under the linguistic observation presented in §2.2, the prior-based metric shares noise-distinguishing behavior with PPL, regardless of the scale of the model or corpus. In particular, (2) we demonstrate that the distance from the central point is a distinguisher of outliers. (3) We also explain the boundary conditions under which this property does not hold.
B.1 Theoretical guarantee of noise detection behavior
B.1.1 The components of PPL and the approximation via prior-based metric
As is well known, PPL consists of the following two components: the likelihood and the prior term.
$\log PPL(d)\propto\sum_{i}\log p(x_{<i}|x_{i})+\sum_{i}\log p(x_{i})$
And in §3.2, we introduced two prior-based metrics:
(1) $\mu_{d}\propto\sum_{i}\log p(x_{i})$
(2) $\sigma_{d}=std[p(x_{i})]\propto\sum_{i}[p(x_{i})-e_{d}]^{2},\ e_{d}=E[p(x_{j})]$
Prior term of PPL
As can be seen, the $\mu_{d}$ is identical to the prior term of PPL. Therefore, the prior-based filter partially approximates PPL.
Likelihood term of PPL
The $\sigma_{d}$ and likelihood share a similarity in that they reflect relationships among tokens rather than averaging the independent properties of individual tokens. However, it is difficult to formally prove a universal connection.
However, under the following assumptions and scenarios, they exhibit similar behavior in filtering noisy data.
Assumptions:
We assume very simplified settings of linguistic observation in §2.2.
(1) Suppose tokens consist of only two categories, F and C, where F represents function words and C represents content words. (2) Within any well-formed document, these two groups must follow a fixed ratio $r^{*}$ (e.g., F:C=4:1).
(3) Actually, data $d$ has a fixed length (e.g., 5), and has an observed ratio $r$. And $r$ has only two possible values: $r^{*}$(well-formed) or $\hat{r}$ (noise).
Example of $r^{*}$ : FFFCF, FFCFF
Example of $\hat{r}$ : FFFFF, FFCCF
(4) If the data of $r^{*}$ takes the big proportion in the corpus, both the empirical average of ratio and the trained model $p$ converge near to $r^{*}$.
(1)–(3) are simplified versions of the linguistic observation in §2.2, and (4) represents the minimal condition required for our method to operate. (Experiment in §3.4.2 shows the minimal proportion is around 10
Property of conditional likelihood
In this simplified setting, conditional likelihood is determined solely by $r$ (e.g., $p(\text{else}\mid C)\ll p(\text{FFFF}\mid C)\approx 1$ ).
Therefore, the distribution of $l_{d}=\sum_{i}\log p(x_{\neq i}|x_{i})$ places data with r* near to 1, while placing $\hat{r}$ near the left tail. And since $r^{*}$ dominates the corpus (assumption 4), the overall distribution of $l_{d}$ becomes right-skewed.
Therefore, in the distribution of $l_{d}$, data with ratio $\hat{r}$ can be identified through the left tails that deviate far from the representative value (e.g., median) of the distribution.
Property of $\sigma_{d}$
For the $\sigma_{d}$, the computation is straightforward. Given a fixed occurrence probabilities (e.g., $p(C)=1/5$ and $p(F)=4/5)$, the $\sigma_{d}$ are distributed around specific value (e.g., $E[\sigma_{d}]=0.26$) where the ratio of data is $r^{*}$. And the samples that deviate far from this point can be distinguished as the one with $\hat{r}$.
Therefore, both the likelihood term and the $\sigma_{d}$ exhibit similar behavior in identifying $\hat{r}$.
Property of $\mu_{d}$
Under the FC assumption, we further explain the role of the prior term of PPL ($\mu_{d}$) in distinguishing $\hat{r}$. Suppose that F and C have fixed occurrence probabilities (e.g., p(F)=4/5,p(C)=1/5p(F)=4/5,\ p(C)=1/5) and appear within each document according to those probabilities. Then, we can observe that the average of their occurrence probabilities tends to concentrate around a specific value (e.g., $E(\mu_{d})=0.68$).
If the observed ratio deviates from this central value, it will move further away from the central value, allowing us to identify $\hat{r}$ accordingly.
In conclusion, under assumptions 1–4, we can guarantee two properties of prior-based metrics. (1) the prior-based metrics ($\mu_{d},\sigma_{d}$) align with all components of PPL in identifying $\hat{r}$. So, we can say the prior-based filter “weakly aligns” with PPL-based filter. (2) the prior-based metric of normal data with valid structures ($r^{*}$) converges around a single central point, allowing outliers ($\hat{r}$) to be distinguished by their distance from this center.
B.1.2 Discrepancy from theoretical guarantee: boundary condition analysis
However, such an approximation required simplified assumptions. Consequently, when these assumptions are violated in a real-world data, a discrepancy arises. The most significant violation case we can identify is the following: even when a document follows the ratio $r^{*}$, it may still contain no meaningful information. This corresponds to cases such as:
(1) Text with an appropriate balance ($r^{*}$) of function and content words, yet are semantically nonsensical. This case is illustrated with Figure 3(a). (2) The data are not noise, nor semantically nonsensical; rather, they are either too difficult or too trivial to be useful for the model. (In this scenario, the main LLM should compute PPL directly, so this case is not applicable to our PPL baseline either.)
In such cases, PPL will play a clearer role, and the prior-based filter will not be able to handle this, and we do not claim this. Prior filter is designed to identify only the noisy data ($\hat{r}$), as we emphasized in the title and §1.
However, we claim the following: (1) The prior can at least provide a clearer signal in identifying $\hat{r}$, as model-predicted likelihood can often be unreliable. (2) In the corpus with a very high proportion of noise (e.g., web crawled corpus), identifying $\hat{r}$ may be more impactful than deliberate reordering among $r^{*}$ data. (3) Still, compatible with PPL-based filter.
B.2 Robustness justification for median-based thresholding
By drawing on the above theoretical arguments, we can justify the robustness of our methodological choices of a median-based threshold. For example, this allows us to answer the question of whether the median-based threshold remains effective when the corpus exhibits a mixed multimodal distribution.
B.2.1 Mixed multimodal distributions
(1) Prior-based filter only measures the grammatical structure via the function-content word ratio. And texts in one specific language tend to share the same grammatical structure, regardless of their style or domain (e.g., news, forum). Thus, prior-filter will generalize to each other.
Dolma is such an example. Although Dolma consists of various types of data (e.g., news, books, symbolic scripts), our method has been shown to perform well across these domains.
(2) Std metric ($\sigma_{d}$) measures the distance between token priors, thus the distance itself may not vary much among other types of language.
Let’s bring the “FC” assumption from §B.1 to here. Let’s assume there are heterogeneous language groups A and B, and each has its function-content words pair, and each group follows the assumption of §B.1. $\{F,C\}\in A,\ \{R,T\}\in B$.
According to the linguistic axiom, we can assume that distinct linguistic system (A, B) shares a similar function-content ratio (e.g., 1:4). Therefore, the central value of std metric for both languages is the same (e.g., $E[\sigma^{A}]=E[\sigma^{B}]=0.26$). Thus, $\sigma_{d}$ for A and B share the same central value, even though they are heterogeneous.
(3) However, in practice, if the amount of B is significantly smaller than A, then there can be an imbalance of the function-content ratio. For example, $p(F)=0.4,p(C)=0.1,p(R)=0.0004,p(T)=0.0001$. In this case, the central value of the distribution of $\sigma^{A}$ and $\sigma^{B}$ will diverge. We suggest two points:
(3-1) We tested an extreme case of this in §3.4.2 with the Chinese-English set, to answer the question “When two entirely different languages are mixed together, to what extent can a single median encompass both languages?” Our experiments indicate that if a minority group is mixed only over 10%, most of the minority language is still considered an inlier when using a single median.
(3-2) If this is still insufficient, we can compute the token prior by mixing in a corpus that shares attributes with the target corpus. For example, if we mix in more data that resembles corpus B and then recalculate the token prior, we might obtain token priors such as $p(F)=0.4$, $p(C)=0.1$, $p(R)=0.4$, and $p(C)=0.1$. In this case, we can bring corpora A and B to the same median value again, even though the actual proportion of B is very small.
This approach corresponds to the method described in §D.1(2) of our paper. For example, when we compute the token prior using a setting where GitHub data are added to Dolma, the resulting token prior enables better identification of GitHub-like data (e.g., symbolic scripts) within Dolma.
Appendix C Generalization to other languages and tokenizers
In this section, we discuss how our method generalizes to languages beyond English and to tokenizers other than the GPT-2 tokenizer.
As we discussed in §B, our method is supposed to work with any language system following two features (1) the language is composed of function words and content words, (2) function and content words have distinct distribution of frequency. This character applies to most natural languages (e.g., Chinese, Japanese, Turkish).
Most languages are composed of function-content words
It is well observed in linguistics (Stromswold, 1994; Croft, 2002) that no matter how compound a language is, it can be separated into function morphemes and content morphemes.
For instance, the Turkish compounding word evlerimizden (‘from our homes’) decomposes into ev (content morpheme) + ler + imiz + den (grammatical morphemes).
These function and content words follow the frequency rule (i.e., function words tend to have higher frequency). For example, in Turkish, the function word “ler” appears across many different nouns (e.g., evler, şehirler, öğrenciler ), resulting in higher term frequency.
The same phenomenon occurs even when dealing with English. For example, a compound word such as childish is typically segmented into child and -ish. The suffix -ish appears across many different content words, and thus behaves similarly to a high-frequency function morpheme.
The tokenizers split words into morphemes based on frequency
Most tokenizers (e.g., BPE, WordPiece, SentencePiece) have tendency to merge characters with high co-occurrence probability into a single token (e.g., “i” + “n” → “in”). This is because most tokenizers, either directly or approximately, follow the Minimum Description Length (MDL) (Hou et al., 2023) objective, which aims to minimize both the vocabulary size and the length of representation. For example, the token “ler” frequently co-occurs with various words, so assigning it as an independent token is more efficient (i.e., covering a wider range of expression with limited vocab) than assigning “şehirler” and “evler” as separate tokens.
Combining together, most of natural languages can be segmented into function and content components, and most tokenizers effectively reflect this distinction, even in morphologically compounding languages.
C.1 Empirical evidence
Table 4: Same setting as Figure 5. Proportion of Turkish data classified as outliers after mixing Turkish and English data at a ratio of $a:100$. Outliers correspond to the top and bottom 5% of $\mu_{d}$.
Size of Turkish data ($a$) Outliers (10%)
1 0.5036
5 0.2778
10 0.2448
20 0.1984
100 0.1337
Other languages
The Chinese–English experiment (§3.4.2) provides evidence for generalization to other languages, such as Chinese.
Suppose that the prior-based filter works only for English but not for Chinese, then our filter will classify all Chinese data as noise. Then, in Figure 5, at the point where Chinese and English are mixed at a 10:100 ratio (at any ratio), nearly 100% of the Chinese data should be classified as outliers. However, the actual results show that most of the Chinese data were classified as inliers, only if this language group takes more than 10% of the corpus. Therefore, the prior-based filter works for Chinese as well.
We also conducted the same experiment on Turkish (Table 4). And the results were consistent with those of the Chinese–English experiment. We provide an explanation below.
Other tokenizers
We provide comparison with two additional tokenizers: LLaMA-3-8B (UTF-8-based, vocab size 128K), and T5-small (SentencePiece-based, vocab size 32K), in Table 5. GPT-2 tokenizer (mainly used in our paper, 50K) has a vocab size of 50K. All of these tokenizers tend to assign function morphemes as individual tokens in accordance with the MDL principle. However, we measure the performance differences among different granularities.
Table 5: Performance of large (1.5B) models on Dolma under different tokenizers. Avg: Avg normalized accuracy, W: World knowledge, C: Commonsense reasoning, L: Language understanding, S: Symbolic problem solving, R: Reading comprehension
Tokenizer Avg W C L S R
GPT2 9.20 9.53 11.27 10.31 11.13 3.79
LLaMA-3 9.39 9.54 11.16 10.78 11.86 3.64
T5-small 8.11 8.59 7.43 7.95 12.32 4.22
The result is reported in ETable 4 (also Table 4 of our paper). All tokenizers showed the advantage of filtering, while T5 exhibited slightly lower performance gain. We propose two possible interpretations: (1) There may exist a threshold of optimal granularity, which lies between 32K and 50K. (2) Using the tokenizer paired with the model may offer stability.
Nevertheless, as demonstrated, priors can be computed even with tokenizers mismatched to the model. If an optimal tokenizer exists, it can be freely adopted at any time.
C.2 Symbolic language
Symbolic languages (e.g., codes) are a more worrisome case, as they are more different from English, and it has not been verified whether linguistic observation of natural language applies to them. Therefore, we conducted experiments on this extreme case (§4.2), and empirically proved our method still generalizes to symbolic language as well. We can expect that it will generalize to other natural languages more similar to English.
Appendix D Preserving minority data